# The role of spatial dimension in the emergence of localised radial patterns from a Turing instability

### v1

@@ 1 Introduction
where $\mathbf{u}\in\mathbb{R}^{2}$, $\Delta:=\partial_{x_{1}}^{2}+\dots+\partial_{x_{n+1}}^{2}$ is the $(n+1)$-dimensional Laplace operator, $\mathbf{D}$ is the diffusion coefficient matrix, $\mathbf{f}$ is a nonlinear function, and $\mu\in\mathbb{R}$ is the bifurcation parameter. Reaction-diffusion systems of the form (1.1) have been found to model numerous physical phenomena, including the growth of desert vegetation [38, 41], the spread of epidemics [3], and nonlinear optics [24]. Beyond specific applications, there are also numerous studies that consider Turing instabilities in $n$ spatial dimensions [36, 2]. Fully localised patterns, consisting of compact patches of spatially-oscillating structure surrounded by a uniform state, have been numerically found to emerge as solutions of (1.1) when the uniform state undergoes a Turing instability [23, 11]. However, our mathematical understanding of the emergence of such patterns remains quite limited when the spatial dimension is larger than one; see [7] for a recent review on the mathematical study of localised patterns.
A notable subclass of fully localised patterns are those which depend purely on the radial coordinate $r:=|\mathbf{x}|$. We refer to such solutions as radial patterns, but they are also called axisymmetric or spherically-symmetric patterns in two and three-dimensions, respectively. Such two-dimensional patterns have been found in vegetation patches and rings [29, 15], chemical reactions [37] and as Kerr solitons [27, 28], while three dimensional localised radial patterns have been observed in Belousov–Zhabotinsky reactions [8] and as cavity bullets in nonlinear optics [34]. When restricting to stationary radial solutions, (1.1) becomes
@@
where the $(n+1)$-dimensional radial Laplace operator takes the form $\Delta_{n}:=\partial^{2}_{r}+\tfrac{n}{r}\partial_{r}$. The spatial dimension $n$ can then be treated as a parameter of (1.2), where one might expect solutions of (1.2) to depend continuously on $n$. Of course, while we can consider $n\in\mathbb{R}$ and find localised solutions of (1.2), these solutions only correspond to localised radial solutions of (1.1) in the case when $n\in\mathbb{N}$. However, the dependence of each solution on the spatial dimension $n$ provides new insight into the connection between two, three, and $n$-dimensional patterns that would be extremely difficult to deduce using Cartesian coordinates.
@@ 2 Main Results
Here we have defined $\mathbf{M}_{1}:=\mathbf{D}(\mu^{*})^{-1}D_{\mathbf{u}}\big{[}\mathbf{f}(\mathbf{u}^{*},\mu^{*})\big{]}$ and introduced $\mathbf{M}_{2}\in\mathbb{R}^{2\times 2}$, $\mathbf{Q}:\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$ and $\mathbf{C}:\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$, where $\mathbf{Q}$ and $\mathbf{C}$ are symmetric bilinear and trilinear maps, respectively. We note that the remainder terms do not affect the subsequent results and so we omit them from our analysis.
With the following hypothesis we make the assumption that (2.1), and hence (1.2), undergoes a nondegenerate Turing instability from the uniform equilibrium $\mathbf{u}=\mathbf{u}^{*}$ at the point $\mu=\mu^{*}$. In what follows $\mathbbm{1}_{m}$ will denote the $m\times m$ identity matrix.
@@
    for some fixed $k_{c}\in\mathbb{R}^{+}$. Furthermore, the eigenvalue $\lambda=-k_{c}^{2}$ of $\mathbf{M}_{1}$ is algebraically double and geometrically simple with generalised eigenvectors $\hat{U}_{0},\hat{U}_{1}\in\mathbb{R}^{2}$, defined so that
@@
The functions $\{J_{0}^{(n)}(r),rJ_{1}^{(n)}(r),Y_{0}^{(n)}(r),rY_{1}^{(n)}(r)\}$ thus form a linearly independent set of solutions to the equation $(\Delta_{n}+1)^{2}u=0$, which corresponds to the linearisation of (2.1) at $\mu=0$. In Lemma 3.1 we use a variation-of-constants formula with this set of linear solutions in order to construct core manifold $\mathcal{W}^{cu}_{-}(\mu)$ on a bounded interval $r\in[0,r_{0}]$.
@@
It is straightforward to prove this hypothesis when $n=2$ as the differential operator becomes the three-dimensional radial Laplacian and one can employ theory from the vast literature on elliptic PDEs (see Lemma 4.3 in [26] and the references therein). However, it is not clear that this result holds for any other values of $n>0$, this hypothesis was eventually proven in the case when $n=1$ by van den Berg et al. in [35] via a computer-assisted proof.
In this work, we present a simplified proof of Hypothesis 2.6 for any choice of $n>0$. Our proof relies on the following proposition which can be proven using standard elliptic PDE theory.
@@
has a positive radial ground state solution $u(\mathbf{x})=Q(|\mathbf{x}|)$ for any $n>0$. In addition, the linearisation of (2.10) about $Q(|\mathbf{x}|)$
@@
Existence of a radial positive ground state follows from standard variational methods for $n>2$ (see for example [20, 21, 33]), and follows directly from [12] for $0<n\leq 2$. The nonexistence of bounded solutions to (2.11) follows from the proof of Lemma 2.1 in [10], where we use the fact that the two quantities $Q(r)$, $Q_{1}(r):=r\mathcal{D}_{\frac{4-n}{2}}Q(r)$ satisfy $\mathcal{L}Q(|\mathbf{x}|)=-2|\mathbf{x}|^{2-n}Q(|\mathbf{x}|)^{3}$ and $\mathcal{L}Q_{1}(|\mathbf{x}|)=-2Q(|\mathbf{x}|)$. ∎
In Lemma 3.4 we show that Proposition 2.7 implies Hypothesis 2.6 and thus guarantees the existence of localised ring and spot B solutions to (2.1) for $40<n<$ 4. We present our key result in the following theorem.
@@
For planar patterns $\nu_{n}=\frac{1}{2}\sqrt{\frac{\pi}{6}}$ and for spherical patterns $\nu_{n}=\frac{\pi}{8}$, which are both in agreement with [22, 26]. While we are unable to derive a closed form expression for $q_{n}$ in the same way as for $\nu_{n}$, we can numerically compute its value for different choices of $n$. We present plots of both $\nu_{n}$ and $q_{n}$ in Figure 2.
@@
where we recall that $\mathbb{O}_{2}$ and $\mathbbm{1}_{2}$ denote the $2\times 2$ zero and identity matrices, respectively.
@@ 3.1 The Core Manifold
We construct the core manifold $\mathcal{W}_{-}^{cu}(\mu)$, for $0<\mu\ll 1$, which contains all small-amplitude solutions to (2.1) that remain bounded as $r\to 0$. This is a local invariant manifold, and so we determine it on a bounded interval $r\in[0,r_{0}]$, for some fixed $r_{0}>0$. We first note that, at the bifurcation point $\mu=0$, the linearisation of (3.1) about $\mathbf{U}=\mathbf{0}$, given by
$\frac{\textnormal{d}}{\textnormal{d}r}\mathbf{V}=\mathcal{A}(r)\mathbf{V},$ (3.2)
has solutions of the form
@@
To see how we arrive at the above solutions for $\mathbf{u}$, we recall that the system $\frac{\textnormal{d}}{\textnormal{d}r}\mathbf{V}=\mathcal{A}(r)\mathbf{V}$ is equivalent to
@@
where $\mathbf{V}=(\mathbf{u},\mathcal{D}_{0}\mathbf{u})^{T}$. Decomposing $\mathbf{u}=u_{1}\hat{U}_{0}+u_{2}\hat{U}_{1}$, we then obtain
@@
We now present a standard result regarding the existence and parametrisation of $\mathcal{W}^{cu}_{-}(\mu)$ on the interval $[0,r_{0}]$, for $0<\mu\ll 1$. In what follows we will use the Landau symbol $\mathcal{O}_{r_{0}}(\cdot)$ with the meaning of the standard Landau symbol $\mathcal{O}(\cdot)$ except that the bounding constants may depend on the value of $r_{0}$.
@@
This statement is proven in a similar way to several previously proven lemmata; see, for example, [22, Lemma 1] or [16, Lemma 4.1]. To do this, we first note that the linear adjoint problem $\frac{\textnormal{d}}{\textnormal{d}r}\mathbf{W}=-\mathcal{A}^{T}(r)\mathbf{W}$ has independent solutions of the form
@@
holds for all $r\in\mathbb{R}$, $i,j\in\{1,2,3,4\}$. For a given $\mathbf{d}=(d_{1},d_{2})\in\mathbb{R}^{2(N+1)}$, we consider the fixed-point equation
@@
on $C([0,r_{0}],\mathbb{R}^{4})$ and prove that (3.10) defines a contraction for sufficiently small $\mathbf{d}$ and $\mu$. Evaluating (3.10) at $r=r_{0}$, we arrive at
@@
$\begin{split}\nu_{n}={}&\frac{1}{2}\left(\frac{\sqrt{\frac{\pi}{2}}}{2^{(\frac{n-1}{2})}\Gamma(\frac{n+1}{2})}\right)^{3}\int_{0}^{\infty}s^{n}(J_{0}^{(n)}(s))^{3}\,\textnormal{d}s,\\ ={}&\frac{\pi\sqrt{\pi}}{4\sqrt{2}}\int_{0}^{\infty}s^{1-(\frac{n-1}{2})}(J_{\frac{n-1}{2}}(s))^{3}\,\textnormal{d}s,\\ ={}&\left(\frac{3}{8}\right)^{\frac{n}{2}}\frac{\pi}{3\Gamma(\frac{n}{2})}.\\ \end{split}$ (3.16)
@@
We conclude this section by noting the following. For sufficiently large $r_{0}$, we can use Table 1 to write (3.7) in terms of its elements $\mathbf{u}(r_{0}),\mathbf{v}(r_{0})\in\mathbb{R}^{2}$, where
$\begin{split}\mathbf{u}(r_{0})&=r_{0}^{-\frac{n}{2}}\left[d_{2}r_{0}\left(1+\mathcal{O}(r_{0}^{-1})\right)\sin\left(y_{n}\right)+d_{1}\left(1+\mathcal{O}(r_{0}^{-1})\right)\cos\left(y_{n}\right)+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|\mathbf{d}|^{2})\right]\hat{U}_{0}\\ &\qquad+2r_{0}^{-\frac{n}{2}}\Big{[}\gamma\,d_{1}^{2}\left(\nu_{n}+\mathcal{O}(r_{0}^{-\frac{n}{2}}+r_{0}^{-1})\right)\sin\left(y_{n}\right)+d_{2}\left(1+\mathcal{O}(r_{0}^{-1})\right)\cos\left(y_{n}\right)\\ &\hskip 233.3125pt+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|d_{2}|^{2}+|d_{1}|^{3})\Big{]}\hat{U}_{1}\\ \mathbf{v}(r_{0})&=r_{0}^{-\frac{n}{2}}\left[d_{2}r_{0}\left(1+\mathcal{O}(r_{0}^{-1})\right)\cos\left(y_{n}\right)-d_{1}\left(1+\mathcal{O}(r_{0}^{-1})\right)\sin\left(y_{n}\right)+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|\mathbf{d}|^{2})\right]\hat{U}_{0}\\ &\qquad+2r_{0}^{-\frac{n}{2}}\Big{[}\gamma\,d_{1}^{2}\left(\nu_{n}+\mathcal{O}(r_{0}^{-\frac{n}{2}}+r_{0}^{-1})\right)\cos\left(y_{n}\right)-d_{2}\left(1+\mathcal{O}(r_{0}^{-1})\right)\sin\left(y_{n}\right)\\ &\hskip 233.3125pt+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|d_{2}|^{2}+|d_{1}|^{3})\Big{]}\hat{U}_{1}\\ \end{split}$ (3.17)
Here we have written $y_{n}:=r_{0}-\frac{n\pi}{4}$, while the $\mathcal{O}_{r_{0}}(\cdot)$ remainders capture the higher order terms when $|\mathbf{d}|$ and $|\mu|$ are taken to be small.

### v3

@@ 1 Introduction
where $\mathbf{u}\in\mathbb{R}^{2}$, $\Delta:=\partial_{x_{1}}^{2}+\dots+\partial_{x_{n+1}}^{2}$ is the $(n+1)$-dimensional Laplace operator, $\mathbf{D}$ is the diffusion coefficient matrix, $\mathbf{f}$ is a nonlinear function, and $\mu\in\mathbb{R}$ is the bifurcation parameter. Reaction-diffusion systems of the form (1.1) have been found to model numerous physical phenomena, including the growth of desert vegetation [43, 46], the spread of epidemics [3], and nonlinear optics [28]. Beyond specific applications, there is also interest in considering Turing instabilities in $n$ spatial dimensions [41, 2]. Fully localised patterns, consisting of compact patches of spatially-oscillating structure surrounded by a uniform state, have been numerically found to emerge as solutions of (1.1) when the uniform state undergoes a Turing instability [27, 11]. However, our mathematical understanding of the emergence of such patterns remains quite limited when the spatial dimension is larger than one; see [6] for a recent review on the mathematical study of localised patterns.
A notable subclass of fully localised patterns are those which depend purely on the radial coordinate $r:=|\mathbf{x}|$. We refer to such solutions as radial patterns, but they are also called axisymmetric or spherically-symmetric patterns in two and three-dimensions, respectively. Such two-dimensional patterns have been found in vegetation patches and rings [33, 18], chemical reactions [42], barrier gas discharge [16], laser processing [23], and as Kerr solitons [31, 32]; see [10] and the references therein for more examples of radial patterns in physical contexts. Three dimensional localised radial patterns are less well known, but have been observed in Belousov–Zhabotinsky reactions [7] and as cavity bullets in nonlinear optics [39]. When restricting to stationary radial solutions, (1.1) becomes
@@
where the $(n+1)$-dimensional radial Laplace operator takes the form $\Delta_{n}:=\partial^{2}_{r}+\tfrac{n}{r}\partial_{r}$. To see this, note that $r=(x^{2}_{1}+x^{2}_{2}+\dots+x_{n+1}^{2})^{\frac{1}{2}}$ and $\partial_{x_{i}}\mathbf{u}(r)=\frac{x_{i}}{r}\mathbf{u}^{\prime}(r)$ for each $i=1,2,\dots,n+1$. The spatial dimension $n$ can then be treated as a parameter of (1.2), where one might expect solutions of (1.2) to depend continuously on $n$. Of course, while we can consider $n\in\mathbb{R}$ and find localised solutions of (1.2), these solutions only correspond to localised radial solutions of (1.1) in the case when $n\in\mathbb{N}$. However, the dependence of each solution on the spatial dimension $n$ provides new insight into the connection between two, three, and $n$-dimensional patterns that would be extremely difficult to deduce using Cartesian coordinates.
@@ 2 Main Results
Here we have defined $\mathbf{M}_{1}:=\mathbf{D}(\mu^{*})^{-1}D_{\mathbf{u}}\big{[}\mathbf{f}(\mathbf{u}^{*},\mu^{*})\big{]}$ and introduced $\mathbf{M}_{2}\in\mathbb{R}^{2\times 2}$, $\mathbf{Q}:\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$ and $\mathbf{C}:\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$, where $\mathbf{Q}$ and $\mathbf{C}$ are symmetric bilinear and trilinear maps, respectively. We remove the remainder terms from the rest of our analysis, as they do not affect the overall results.
The following hypothesis establishes that (2.1), and thus (1.2), undergoes a nondegenerate Turing instability from the uniform equilibrium $\mathbf{u}=\mathbf{u}^{*}$ at the point $\mu=\mu^{*}$.
@@
    for some fixed $k_{c}\in\mathbb{R}^{+}$, where $\mathbbm{1}_{2}$ is the two-dimensional identity matrix. Furthermore, the eigenvalue $\lambda=-k_{c}^{2}$ of $\mathbf{M}_{1}$ is algebraically double and geometrically simple with generalised eigenvectors $\hat{U}_{0},\hat{U}_{1}\in\mathbb{R}^{2}$, defined so that
@@
The functions $\{J_{0}^{(n)}(r),rJ_{1}^{(n)}(r),Y_{0}^{(n)}(r),rY_{1}^{(n)}(r)\}$ thus form a linearly independent set of solutions to the equation $(\Delta_{n}+1)^{2}u=0$, which corresponds to the linearisation of (2.1) at $\mu=0$. In Lemma 3.1 we use a variation-of-constants formula with this set of linear solutions in order to construct core manifold $\mathcal{W}^{cu}_{-}(\mu)$ on a bounded interval $r\in[0,r_{0}]$. We emphasise that our use of this set of functions is not an assumption, but rather an enforced choice based on the structure of the linear operator $(\Delta_{n}-\mathbf{M}_{1})\mathbf{u}=\mathbf{0}$.
@@
It is straightforward to prove this hypothesis when $n=2$ as the differential operator becomes the three-dimensional radial Laplacian and one can employ theory from the vast literature on elliptic PDEs (see Lemma 4.3 in [30] and the references therein). However, it is not clear that this result holds for any other values of $n>0$, this hypothesis was eventually proven in the case when $n=1$ by van den Berg et al. in [40] via a computer-assisted proof. Based on the expected asymptotic scaling $|\mathbf{u}|\sim\mu^{\frac{4-n}{8}}$ for spot B solutions in [30], as well as a similar prediction of $|\mathbf{u}|\sim\mu^{\frac{4-n}{4}}$ for ring solutions, we only consider solutions to (2.8) for $40<n<$ 4. We will observe that this is required in order to ensure certain terms are asymptotically small in our subsequent analysis.
In this work, we present a simplified version of Hypothesis 2.6 which can be proven for any choice of $30<n<$ 3. Our proof relies on the following proposition which can be proven using standard elliptic PDE theory.
@@
has a positive radial ground state solution $u(\mathbf{x})=Q(|\mathbf{x}|)$ for any $30<n<$ 3. In addition, the linearisation of (2.10) about $Q(|\mathbf{x}|)$
@@
Existence of a radial positive ground state follows from standard variational methods, see for example [24, 25, 37]. In particular, existence was proven for $32<n<$ 3 in [13], and follows directly from [12] for $0<n\leq 2$. The nonexistence of bounded solutions to (2.11) follows from the proof of Lemma 2.1 in [9], where we use the fact that the two quantities $Q(r)$, $Q_{1}(r):=r\mathcal{D}_{\frac{4-n}{2}}Q(r)$ satisfy $\mathcal{L}Q(|\mathbf{x}|)=-2|\mathbf{x}|^{2-n}Q(|\mathbf{x}|)^{3}$ and $\mathcal{L}Q_{1}(|\mathbf{x}|)=-2Q(|\mathbf{x}|)$. ∎
In Lemma 3.4 we show that Proposition 2.7 implies Hypothesis 2.6 and thus guarantees the existence of localised ring and spot B solutions to (2.1) for $30<n<$ 3. We also show the existence of localised ring solutions for $43\leq n<$ 4, subject to the following hypothesis.
Hypothesis 2.8.
Proposition 2.7 also holds for any value $43\leq n<$ 4.
The standard techniques used to prove the existence of $H^{1}(\mathbb{R}^{3})$ (and consequently $C^{2}(\mathbb{R}^{3})$) radial solutions in Proposition 2.7 break down when $n>3$. In fact, one can prove the nonexistence of $H^{1}(\mathbb{R}^{3})$ solutions to (2.10) in this regime. It may still be possible to obtain solutions in a different function space that guarantees the existence of localised ring solutions for $43\leq n<$ 4, however we leave this question to future study and note the following. For the remainder of this work, the restriction $40<n<$ 4 for the existence of ring and spot B solutions is to be understood as either $30<n<$ 3 or $43\leq n<$ 4 subject to Hypothesis 2.8.
We present our key result in the following theorem.
@@
For planar patterns $\nu_{n}=\frac{1}{2}\sqrt{\frac{\pi}{6}}$ and for spherical patterns $\nu_{n}=\frac{\pi}{8}$, which are both in agreement with [26, 30]. The implicit curve (2.16) implies that, for a fixed $0<\gamma\ll 1$, the fold point $\mu=\mu_{*}$ with $0<\mu_{*}\ll 1$ must increase as $n$ increases. This is in agreement with Figure 1, although we note that the parameter values in Figure 1 are far from the values where our analysis holds. While we are unable to derive a closed form expression for $q_{n}$ in the same way as for $\nu_{n}$, we can numerically compute its value for different choices of $n$ by solving (2.9) via finite difference techniques. We present plots of both $\nu_{n}$ and $q_{n}$ in Figure 2.
@@
where we note that $\mathbb{O}_{2}$ and $\mathbbm{1}_{2}$ are the zero and identity matrices in $\mathbb{R}^{2\times 2}$, respectively.
@@ 3.1 The Core Manifold
We begin by characterising the core manifold $\mathcal{W}_{-}^{cu}(\mu)$, which forms a local invariant manifold containing all small-amplitude bounded solutions to (2.1) on the interval $r\in[0,r_{0}]$, for some large $r_{0}>0$. At the bifurcation point $\mu=0$, the linearisation of (3.1) has solutions of the form
@@
These solutions follow by first noting that the system $\frac{\textnormal{d}}{\textnormal{d}r}\mathbf{V}=\mathcal{A}(r)\mathbf{V}$ is equivalent to
@@
with $\mathbf{V}=(\mathbf{u},\mathcal{D}_{0}\mathbf{u})^{T}$. By taking the eigenbasis decomposition $\mathbf{u}=u_{1}\hat{U}_{0}+u_{2}\hat{U}_{1}$, we then obtain
@@
We now present a standard result regarding the existence and parametrisation of $\mathcal{W}^{cu}_{-}(\mu)$ on the interval $[0,r_{0}]$, for $0<\mu\ll 1$. We use the notation $\mathcal{O}_{r_{0}}(\cdot)$ to denote the standard Landau symbol $\mathcal{O}(\cdot)$ with bounding constants that depend on the value of $r_{0}$. Additionally, we define $\mathcal{P}^{cu}_{-}(r_{0})$ to be a projection onto the subspace spanned by bounded (as $r\to 0$) solutions $\{\mathbf{V}_{1}(r_{0}),\mathbf{V}_{2}(r_{0})\}$, evaluated at the point $r=r_{0}$.
@@
This statement is proven in a similar way to several previously proven lemmata; see, for example, [26, Lemma 1] or [19, Lemma 4.1]. We first note that the linear adjoint problem $\frac{\textnormal{d}}{\textnormal{d}r}\mathbf{W}=-\mathcal{A}^{T}(r)\mathbf{W}$ is equivalent to $(\mathcal{D}_{n}\mathcal{D}_{0}-\mathbf{M}_{1}^{T})\mathbf{w}=\mathbf{0}$, where $\mathbf{W}=(-r^{n}\mathcal{D}_{0}\mathbf{w},r^{n}\mathbf{w})^{T}$, and thus has independent solutions of the form
@@
holds for all $r>0$, $i,j\in\{1,2,3,4\}$. The orthonormal property follows in part from the identity
$r^{n}[J_{1}^{(n)}(r)Y_{0}^{(n)}(r)-J_{0}^{(n)}(r)Y_{1}^{(n)}(r)]=\frac{2^{n}\Gamma(\tfrac{n+1}{2})^{2}}{\pi}.$
For a given $\mathbf{d}=(d_{1},d_{2})\in\mathbb{R}^{2(N+1)}$, we consider the fixed-point equation
@@
on $C([0,r_{0}],\mathbb{R}^{4})$, whose solutions define all solutions to (3.1) that remain bounded on $[0,r_{0}]$. After proving that (3.9) defines a contraction for sufficiently small $\mathbf{d}$ and $\mu$ and evaluating at $r=r_{0}$, we arrive at
@@
$\begin{split}\nu_{n}={}&\frac{\pi\sqrt{\pi}}{4\sqrt{2}}\int_{0}^{\infty}s^{1-(\frac{n-1}{2})}(J_{\frac{n-1}{2}}(s))^{3}\,\textnormal{d}s=\left(\frac{3}{8}\right)^{\frac{n}{2}}\frac{\pi}{3\Gamma(\frac{n}{2})}.\\ \end{split}$ (3.15)
@@
For sufficiently large values of $r_{0}$, Table 1 allows us to write down the following leading-order expansion for (3.6)
$\begin{split}\mathbf{u}(r_{0})&=r_{0}^{-\frac{n}{2}}\left[d_{2}r_{0}\left(1+\mathcal{O}(r_{0}^{-1})\right)\sin\left(y_{n}\right)+d_{1}\left(1+\mathcal{O}(r_{0}^{-1})\right)\cos\left(y_{n}\right)+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|\mathbf{d}|^{2})\right]\hat{U}_{0}\\ &\qquad+2r_{0}^{-\frac{n}{2}}\Big{[}\gamma\,d_{1}^{2}\left(\nu_{n}+\mathcal{O}(r_{0}^{-\frac{n}{2}}+r_{0}^{-1})\right)\sin\left(y_{n}\right)+d_{2}\left(1+\mathcal{O}(r_{0}^{-1})\right)\cos\left(y_{n}\right)\\ &\hskip 233.3125pt+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|d_{2}|^{2}+|d_{1}|^{3})\Big{]}\hat{U}_{1},\\ \mathbf{v}(r_{0})&=r_{0}^{-\frac{n}{2}}\left[d_{2}r_{0}\left(1+\mathcal{O}(r_{0}^{-1})\right)\cos\left(y_{n}\right)-d_{1}\left(1+\mathcal{O}(r_{0}^{-1})\right)\sin\left(y_{n}\right)+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|\mathbf{d}|^{2})\right]\hat{U}_{0}\\ &\qquad+2r_{0}^{-\frac{n}{2}}\Big{[}\gamma\,d_{1}^{2}\left(\nu_{n}+\mathcal{O}(r_{0}^{-\frac{n}{2}}+r_{0}^{-1})\right)\cos\left(y_{n}\right)-d_{2}\left(1+\mathcal{O}(r_{0}^{-1})\right)\sin\left(y_{n}\right)\\ &\hskip 233.3125pt+\mathcal{O}_{r_{0}}(|\mu||\mathbf{d}|+|d_{2}|^{2}+|d_{1}|^{3})\Big{]}\hat{U}_{1}.\\ \end{split}$ (3.16)
Here we have defined $y_{n}:=r_{0}-\frac{n\pi}{4}$, and the remainder terms $\mathcal{O}_{r_{0}}(\cdot)$ capture the higher order terms for small values of $|\mathbf{d}|$ and $|\mu|$.